A Complete Bibliography of Journal of Data and Information Quality (JDIQ)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 October 2025 Version 1.42

Title word cross-reference

2017 [?]. N [?].

-gram [?]. -November [?].

19 [?, ?, ?, ?, ?].

3D [?]. 3D-DaVa [?].

Abnormal [?, ?, ?]. Academia [?]. Academic [?]. Acceptance [?]. Access [?]. accidents [?]. Accuracy [?, ?, ?, ?, ?]. Achieving [?]. ACM [?, ?]. Across [?, ?]. Actions [?, ?]. Active [?, ?]. Activity [?]. Actors [?]. Adaptation [?]. Adapting [?]. Adaptive [?, ?, ?, ?, ?]. Address [?]. Addressing [?, ?]. Administrations [?]. Advanced [?]. Adversarial [?]. AED [?]. Affecting [?]. Age [?]. Agent [?]. Agent-Based [?]. Agents [?]. aggregate [?]. Aggregation [?]. AI [?, ?, ?, ?, ?, ?, ?, ?]. AI-generated [?]. Algorithm [?, ?, ?, ?]. Algorithmic [?]. Algorithms [?, ?, ?]. Alignment [?]. alive [?]. Ambient [?]. Amounts [?]. Amrapali [?].

Analyses [?]. Analysis [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Analytic [?]. Analytics [?, ?, ?, ?, ?, ?, ?, ?]. Analyzing [?]. Anatomy [?]. annealing [?]. Annotation [?, ?]. Annotations [?]. Anomaly [?, ?]. Anomaly-based [?]. Anonymization [?]. Anserini [?]. Answering [?, ?]. Ante [?]. Application [?, ?, ?, ?]. Applications [?, ?, ?, ?]. Approach [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Architecture [?]. Archives [?]. Archiving [?]. art [?]. Articles [?]. Artificial [?, ?]. Aspect [?]. Aspect-Level [?]. Aspects [?]. Aspell [?]. Assess [?]. Assessing [?, ?, ?, ?]. Assessment [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Assets [?]. Assisted [?, ?, ?]. Association [?]. Assurance [?]. Attempts [?]. Attention [?, ?]. Attribute [?]. Attributes [?]. Audio [?]. Audio-based [?]. Augmentation [?]. Augmenting [?]. Authentication [?, ?]. Automate [?]. Automated [?, ?, ?, ?, ?]. Automatic [?, ?, ?, ?, ?]. Automating [?]. AutoML [?]. Autonomous [?]. Aware [?, ?, ?, ?, ?, ?, ?, ?, ?]. Awareness [?].

Balance [?]. Banking [?]. Base [?, ?]. Based [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Baselines [?]. Bases [?, ?]. Bayesian [?, ?, ?]. Be [?]. Beat [?]. Being [?]. Believing [?]. Benchmark [?]. Benchmarking [?, ?]. Between [?, ?, ?]. Beyond [?]. Bias [?, ?, ?, ?]. Biased [?]. Biases [?, ?]. Big [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. BigCo [?]. BIGQA [?]. Billion [?]. Biological [?]. Biomedical [?]. Biometrics [?, ?]. Bipartite [?]. Bitrate [?]. Black [?]. BLAST2 [?]. Blockchain [?, ?]. Blocking [?]. Book [?]. Breach [?]. Bridging [?, ?]. Briefs [?]. Broader [?]. BUNNI [?]. Business [?].

CAD [?]. Call [?, ?]. Campaign [?]. Campaigns [?]. cancer [?]. Capabilities [?]. Car [?]. Care [?, ?]. Case [?, ?, ?, ?, ?, ?, ?, ?, ?]. Cases [?]. Catalog [?]. Categorical [?, ?]. Categories [?]. Categorization [?]. centric [?]. Challenge [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Challenges [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Challenging [?]. Character [?]. Characteristics [?]. Characterization [?]. Characterizing [?]. Checkers [?]. Checking [?, ?, ?, ?]. chest [?]. Chief [?, ?, ?]. Choice [?]. chronic [?]. Cities [?, ?]. Cities-Framework [?]. Claim [?, ?]. Classification [?, ?, ?, ?, ?, ?, ?, ?]. Classify [?]. Classifying [?]. Classroom [?]. Clean [?]. Cleaning [?, ?, ?, ?, ?, ?]. Cleenex [?]. Clickbait [?]. Client [?]. Client-side [?]. Clinical [?]. Cloud [?, ?, ?]. Cluster [?, ?, ?]. Cluster-Based [?]. Cluster-then-label [?]. Clustering [?, ?, ?]. Clusters [?]. Coding [?]. Cold [?]. Cold-start [?]. Collaborative [?, ?]. Collection [?, ?]. Collections [?]. Combating [?]. Combinations [?]. Combining [?, ?, ?, ?, ?]. Come [?]. Comments [?, ?]. Common [?]. Comparative [?, ?]. Comparison [?, ?, ?]. Compensated [?]. Competitive [?]. Complement [?]. Complete [?]. Completeness [?]. Completing [?].

Complex [?]. Complexity [?]. Compliance [?, ?]. Components [?]. Compound [?]. Compression [?, ?]. Compromising [?]. Computation [?]. Computational [?]. computed [?]. Computer [?]. Computing [?, ?, ?]. Computing-Assisted [?]. Conceptual [?]. Concurrent [?]. Conditional [?]. Conference [?]. Confidence [?]. Conflict [?]. Connected [?, ?]. Connecting [?]. Connection [?]. Connectivity [?]. Consistency [?, ?]. Constraint [?]. Constraints [?, ?, ?]. Construct [?]. Construction [?]. Constructive [?]. Consumer [?]. Content [?, ?, ?, ?, ?, ?, ?, ?, ?]. Content-Aware [?]. Content-based [?]. Context [?, ?, ?, ?, ?, ?, ?]. Context-aware [?]. Context-Driven [?]. Contextual [?, ?, ?]. Contingency [?]. Control [?, ?]. controlled [?]. Controlling [?]. Conversations [?]. Convoluted [?]. Cooperation [?]. Coreference [?]. Corpora [?]. Corpus [?, ?]. Correctness [?]. Corrigendum [?]. Cost [?]. Cost-Effective [?]. costs [?]. counterfeits [?]. Course [?]. Covariates [?]. Coverage [?]. Coverage-based [?]. COVID [?, ?, ?, ?, ?]. COVID-19 [?, ?, ?, ?, ?]. Creating [?]. Creation [?, ?]. Credibility [?, ?]. Criteria [?, ?]. Critical [?, ?]. Criticality [?]. CRM [?]. Cross [?]. Cross-Modal [?]. Crowd [?]. Crowd-Assisted [?]. Crowdsourced [?]. Cryptocurrency [?]. Cumulative [?]. Curated [?]. Curation [?, ?, ?, ?]. Currency [?, ?]. Curtailment [?]. Customer [?, ?]. Customization [?]. Cyber [?]. Cyber-Physical [?]. Cyberbullying [?]. Cycle [?].

Data [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Data [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. data- [?]. Data-centric [?]. Database [?, ?, ?, ?, ?]. Databases [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Dataset [?, ?]. Datasets [?, ?, ?, ?]. DaVa [?]. DB [?]. Debugging [?]. Deception [?, ?]. Deceptive [?, ?]. Decision [?, ?]. Decisions [?, ?, ?]. Declarative [?]. Dedicated [?]. Deduplication [?, ?]. Deep [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Definition [?]. Definitions [?]. Degradation [?]. Delivering [?]. Dependable [?]. Dependencies [?, ?, ?, ?, ?]. Dependency [?]. Descriptors [?]. Design [?, ?, ?, ?, ?, ?, ?, ?, ?]. Designing [?]. Detail [?]. Detecting [?, ?, ?, ?, ?]. Detection [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Deterring [?]. Develop [?]. Developing [?]. Development [?]. Device [?]. Devices [?]. Diagnosis [?]. Diagnosis-Procedure [?]. diagnostic [?]. dictation [?]. Dictionaries [?]. Different [?]. Differentiating [?]. Diffusion [?]. Digital [?, ?, ?, ?, ?, ?, ?]. Dimensional [?]. Dimensions [?, ?, ?]. Directions [?]. Dirty [?]. Disaggregation [?]. Disambiguation [?]. Discourse [?]. Discover [?, ?]. Discovering [?, ?]. Discovery [?, ?, ?, ?]. Discriminant [?]. Discussion [?, ?]. Disease [?, ?]. Disinformation [?, ?]. Distance [?]. Distributed [?, ?]. Distribution [?]. Diversity [?]. Document [?, ?, ?, ?].

Documentation [?]. Does [?]. Domain [?, ?]. Domain-Independent [?]. Domain-Specific [?]. Domains [?, ?]. Driven [?, ?, ?, ?, ?, ?]. Drug [?]. Dual [?]. due [?]. Duplicate [?, ?, ?]. Duplicates [?]. During [?, ?]. DUTNG [?]. Dynamic [?, ?, ?, ?]. Dynamically [?].

E-FAIR-DB [?]. E-learning [?]. Ebola [?]. Ecosystem [?]. Ecosystems [?]. Edge [?, ?]. Edge/Cloud [?]. Edges [?]. Editor [?, ?, ?, ?]. Editor-in-Chief [?, ?, ?]. Editorial [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Editors [?]. Education [?]. Educational [?]. EEG [?]. EEG-Mechatronic [?]. Effective [?]. Effectiveness [?]. Effects [?, ?, ?, ?, ?]. Efficient [?, ?]. Effort [?]. electronic [?, ?]. ELF [?]. Embedding [?, ?]. Emergency [?]. Emerging [?]. Employee [?]. Employing [?]. enabled [?]. Enabling [?]. Encountered [?]. Encryption [?]. Encryption-Compression-based [?]. Engineering [?, ?]. Enhance [?]. Enhanced [?]. Enhancing [?, ?, ?, ?]. Enrichment [?]. Ensemble [?]. Ensure [?]. Ensuring [?, ?]. Enterprises [?]. Entities [?]. Entity [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. entry [?]. Environment [?, ?]. Environments [?, ?, ?, ?, ?]. Epidemiological [?]. Equity [?, ?]. Era [?]. ERP [?]. ERP-enabled [?]. Error [?]. Errors [?, ?]. estate [?]. Estimating [?, ?, ?]. Estimation [?, ?]. eTextbook [?]. Ethical [?, ?]. Ethics [?, ?]. Evaluate [?, ?, ?]. Evaluating [?, ?]. Evaluation [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Evaluation-as-a-Service [?]. Event [?, ?, ?]. Evidence [?]. Evolution [?]. Evolutionary [?]. Ex [?]. Examples [?, ?]. Exchange [?]. Exemplar [?]. ExpanDrogram [?]. Expectations [?]. Experience [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Experiment [?]. Experimental [?, ?, ?]. Expert [?, ?]. Explain [?]. Explainability [?]. Explainable [?]. Explaining [?]. Exploratory [?, ?, ?]. Expressive [?]. Extend [?]. Extended [?]. Extraction [?, ?, ?, ?].

Face [?, ?]. Face-to-Face [?]. Faces [?]. Facets [?]. Fact [?, ?]. Fact-Checking [?]. Failure [?]. FAIR [?, ?]. Fairness [?, ?, ?, ?]. Fairness-aware [?]. Fake [?, ?]. False [?, ?, ?]. Family [?]. Farewell [?, ?]. Feature [?, ?]. Features [?, ?]. Federated [?]. Feedback [?]. Few [?, ?]. Few-shot [?]. Field [?, ?]. File [?]. filling [?]. Financial [?, ?, ?]. Find [?]. Fire [?]. Firm [?, ?]. Firm-level [?]. Firm-Specific [?]. Fiscal [?]. Fit [?]. Fixing [?]. Flow [?, ?]. flows [?]. Footwear [?]. Forecasting [?]. Forensic [?]. Forest [?]. Foreword [?]. form [?]. Format [?]. Forming [?]. Framework [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Fraud [?]. Frequency [?]. Frisk [?]. Frontier [?]. Functional [?, ?, ?]. Fusion [?]. Fusion-based [?]. Future [?, ?].

game [?]. game-theoretic [?]. Gap [?]. Gathered [?]. Gender [?]. General [?, ?]. General-Purpose [?]. Generalizable [?]. Generalize [?]. Generated [?, ?, ?, ?]. Generation [?, ?]. Generative [?]. Geocoding [?]. German [?, ?]. Getting [?]. GIGO [?]. Glitches [?]. Global [?]. Gold [?, ?]. Good [?]. Governance [?, ?, ?, ?]. Government [?]. Graduate [?]. Graduate-Level [?]. gram [?]. Graph [?, ?, ?, ?, ?]. Graph-Based [?]. Graphs [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Growth [?]. Guest [?]. Guided [?].

Handling [?]. Harvesting [?]. health [?, ?]. Healthcare [?, ?, ?, ?, ?, ?]. Help [?]. Heterogeneous [?, ?]. High [?, ?, ?, ?, ?]. High-Dimensional [?]. High-Precision [?]. High-Quality [?, ?, ?]. Historical [?, ?]. Hoaxes [?]. Human [?, ?, ?, ?, ?, ?, ?]. Human-in-the-Loop [?]. Humans [?]. Hunspell [?]. Hunting [?]. Hybrid [?, ?, ?, ?].

Ideas [?]. Identification [?]. Identifying [?, ?]. II [?]. Image [?, ?]. Images [?, ?]. Imaging [?]. Impact [?, ?, ?, ?]. Implementation [?, ?, ?, ?]. implications [?]. Important [?]. Imprecise [?]. Improve [?, ?, ?]. Improved [?]. Improvement [?]. Improving [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Inaugural [?]. Incentive [?]. Incorporating [?]. Incremental [?]. Incubating [?]. Independent [?]. Indeterministic [?]. Indexing [?]. Indicators [?]. Industrial [?]. Industry [?, ?]. Industry-Scale [?]. Inference [?]. InfoClean [?]. Information [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Information-Based [?]. Informed [?]. Infrastructure [?]. Infrastructures [?]. Inpainting [?]. Insights [?, ?]. Instance [?]. Instances [?]. Instrument [?]. Insurance [?]. Integral [?]. Integrated [?]. Integration [?, ?, ?, ?, ?, ?, ?, ?]. Intelligence [?, ?, ?, ?]. Intelligent [?]. Interaction [?]. Interactions [?, ?]. Interactive [?, ?, ?]. Interface [?]. Intermediation [?]. Internet [?, ?]. interoperability [?]. Interoperable [?]. Introduction [?, ?, ?, ?, ?, ?]. Intrusion [?]. Inventory [?]. Involvement [?]. IoMT [?]. IoT [?, ?, ?]. IoT/ML [?]. Ireland [?]. Isolated [?]. Issue [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Issues [?, ?, ?, ?, ?]. Italian [?]. Iterative [?].

January [?]. JDIQ [?, ?, ?, ?, ?]. Joint [?]. Jointly [?]. Journal [?, ?]. Judging [?]. Judgments [?]. June [?]. Just [?].

Knowledge [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Knowledge-Driven [?].

Lab [?]. Label [?, ?, ?]. Labeling [?, ?]. Lack [?]. Language [?, ?, ?]. Languages [?, ?]. Large [?, ?, ?, ?, ?, ?, ?]. Large-scale [?, ?]. Law [?]. Learned [?]. Learner [?]. Learning [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Learning-based [?]. Lessons [?]. Letter [?]. Level [?, ?, ?, ?]. Leverage [?]. Libraries [?, ?, ?]. License [?]. Life [?]. Lightweight [?]. Lineage [?]. Linear [?]. Link [?]. Linkage [?, ?]. Linked [?, ?]. Linking [?]. Links [?]. Literature [?, ?]. Live [?]. Live-Virtual-Constructive [?]. Living [?, ?]. LLM [?, ?]. LLM-based [?]. Location [?]. Log [?]. Logs [?]. Long [?]. Long-Term [?]. Loop [?, ?]. Loose [?]. Low [?]. Low-Quality [?]. Lucene [?]. lung [?].

Machine [?, ?, ?, ?, ?, ?, ?, ?, ?]. Maintenance [?]. Making [?, ?]. Malware [?]. Management [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Management-BigCo [?]. Managing [?]. Manufacturer [?]. manufacturing [?]. Many [?]. Market [?, ?, ?]. Matching [?, ?, ?, ?, ?]. Matter [?]. May [?]. Measure [?]. Measurement [?, ?]. Measures [?, ?]. Measuring [?]. Mechanism [?]. Mechatronic [?]. Media [?, ?, ?, ?]. Media-Insights [?]. Medical [?]. meets [?]. Memes [?]. Memory [?]. Metadata [?, ?, ?, ?]. Method [?, ?, ?, ?, ?, ?]. Methodology [?, ?, ?]. Methods [?, ?, ?, ?]. Metric [?, ?]. Metrics [?, ?, ?, ?, ?, ?, ?]. Metrics-driven [?]. mHealth [?]. Mind [?]. minded [?]. Mine [?]. Mining [?, ?, ?, ?, ?, ?, ?]. Misinformation [?, ?, ?, ?]. Missing [?, ?, ?, ?, ?]. Mixed [?]. ML [?, ?]. Mobile [?, ?, ?]. Modal [?]. Model [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Modeling [?, ?, ?]. Modelling [?]. Models [?, ?, ?, ?, ?, ?, ?]. Multi [?, ?]. Multi-Criteria [?, ?]. Multicriteria [?]. Multidimensional [?, ?, ?]. Multifactor [?]. Multimedia [?, ?, ?, ?, ?]. Multimodal [?, ?, ?]. Multimodality [?]. Multimode [?]. Multivariate [?].

Name [?]. Names [?]. Navigation [?]. Need [?]. Negative [?]. Neighborhood [?]. Network [?, ?, ?, ?]. Networks [?, ?, ?]. Neural [?, ?]. Neutrality [?]. News [?, ?]. Nodes [?]. Noise [?]. Noisy [?, ?]. Nondiscrimination [?]. Nondiscrimination-aware [?]. Normal [?]. Normalization [?]. Notes [?]. Novel [?, ?, ?, ?, ?, ?]. November [?]. Numbers [?]. NYPD [?].

Object [?, ?]. Observability [?]. Observatories [?]. Offensive [?]. One [?]. Online [?, ?, ?, ?, ?]. Ontological [?]. Ontologies [?, ?]. Ontology [?, ?, ?, ?, ?]. Ontology-based [?]. Open [?, ?, ?, ?, ?, ?, ?, ?]. OpenSearch [?]. Operations [?]. Opinion [?]. Opinions [?]. Opportunities [?, ?, ?, ?]. Optimal [?]. order [?]. Origins [?]. Other [?, ?]. Outbreak [?]. Outlier [?]. Outlook [?]. Output [?]. Overheard [?]. Overview [?, ?, ?].

Page [?]. Pairwise [?]. Panama [?]. Panels [?]. Panther [?]. Paper [?, ?, ?, ?]. Papers [?, ?]. Part [?, ?]. Partial [?]. Partial-order [?]. Patient [?, ?]. patient-level [?]. Patterns [?, ?]. Percentages [?]. Perceptions [?]. Performance [?, ?, ?, ?]. Periocular [?]. Personal [?, ?]. perspective [?]. Physical [?]. Pipeline [?]. Pipelines [?]. Platforms [?, ?]. Point [?]. Poisoning [?]. Policies [?]. Policy [?]. Policymakers [?]. Pooling [?]. Positive [?]. Post [?]. Post-Truth [?]. PoWareMatch [?]. Practical [?, ?]. Practice [?, ?, ?]. Precision [?]. Predict [?]. Predicting [?]. Prediction [?, ?, ?, ?]. Predictive [?]. Preference [?, ?]. Premature [?]. Preparation [?, ?, ?]. Prepare [?]. Preprocessing [?]. Presentation [?]. preservation [?]. Pressures [?]. Primary [?, ?]. Principles [?].

Privacy [?, ?, ?]. Private [?, ?]. Probabilistically [?]. Probabilities [?]. Probability [?]. Probability-based [?]. Probably [?]. Problem [?, ?]. problems [?]. Procedure [?, ?, ?]. Process [?, ?, ?, ?, ?]. Process-Data [?]. Process-driven [?]. Processes [?]. Processing [?, ?]. Product [?, ?]. Profiled [?]. Profiles [?]. Profiling [?]. Program [?, ?]. programs [?]. project [?]. Propagation [?]. Property [?]. Protect [?]. Protecting [?]. Protection [?]. Provenance [?, ?, ?, ?, ?]. Provenance-Aware [?]. Public [?, ?]. Publications [?, ?]. Pure [?]. Purpose [?].

QDflows [?]. QUAL [?]. Quality [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Quality [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Quality-Aware [?, ?, ?, ?]. quantifying [?]. Quantitative [?, ?]. Queries [?, ?]. Query [?, ?, ?]. Questions [?]. Queue [?]. Quick [?].

randomized [?]. Randomness [?]. Ranking [?]. RDF [?, ?, ?]. Reach [?]. Reader [?]. Readiness [?, ?]. Reading [?]. Real [?, ?, ?]. Real-estate [?]. Real-Time [?]. Recommendations [?]. Recommender [?, ?]. Record [?, ?, ?, ?, ?]. Record-Linkage [?]. records [?]. Regulation [?]. Regulatory [?]. Relational [?, ?]. Relational-to-RDF [?]. Relationship [?]. Relationship-Based [?]. Relative [?]. Relevance [?]. Reliability [?]. Relief [?]. Remote [?]. Repair [?, ?]. Repairing [?, ?, ?]. Repairs [?, ?]. Replacing [?]. Report [?, ?, ?]. Reporting [?]. Reports [?]. Representation [?, ?, ?]. Representing [?]. Reproduce [?, ?]. Reproducibility [?, ?, ?, ?]. Reproducible [?, ?]. Repurposing [?, ?]. Reputation [?]. Requirements [?, ?, ?, ?]. Research [?, ?, ?, ?, ?, ?, ?]. Resilience [?]. Resolution [?, ?, ?, ?, ?, ?, ?, ?]. Response [?]. Responsible [?, ?, ?]. Results [?, ?, ?]. Retention [?]. Retrieval [?, ?, ?, ?, ?, ?, ?]. Retrospect [?]. Reuse [?]. Review [?, ?, ?, ?, ?, ?, ?]. Revision [?]. Revisiting [?]. Rid [?]. Ring [?]. Risk [?, ?, ?, ?, ?]. Robust [?]. Robustness [?]. Role [?]. Roles [?]. Rule [?]. Rule-driven [?]. RuleHub [?]. Rules [?, ?, ?]. Rumors [?].

Safety [?]. Samples [?]. Sampling [?]. Satisficing [?]. satisfying [?]. Scalable [?]. Scale [?, ?, ?]. Scaling [?]. Scenario [?]. Schema [?, ?]. Scheme [?]. Scholarly [?]. Science [?, ?, ?]. Sciences [?]. Scientific [?]. Scoping [?]. Scoring [?, ?]. Screen [?]. Search [?, ?, ?, ?]. Sector [?]. Securing [?]. Security [?, ?, ?, ?, ?]. Seeing [?]. Segmentation [?, ?]. Select [?]. Selection [?, ?, ?, ?, ?, ?]. Self [?]. Self-Identifying [?]. Semantic [?, ?]. Semantics [?]. Sensing [?]. Sensitive [?]. Sensor [?]. Sensors [?]. Sentiment [?, ?]. Sequence [?, ?, ?]. Series [?, ?, ?, ?]. Serves [?]. Service [?, ?]. Services [?, ?, ?, ?, ?]. Session [?]. Sessions [?]. settings [?]. Shape [?]. Shared [?]. Sharing [?]. Shenanigans [?]. shops [?]. Short [?]. Short-term [?]. shot [?]. Should [?]. Shrinkage [?].

Siamese [?]. side [?]. Signature [?, ?]. Significant [?]. Similarity [?, ?]. Simulate [?]. Size [?]. Smart [?, ?, ?]. SNE [?]. Social [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Social-minded [?]. Societies [?]. Society [?]. Soft [?, ?]. Software [?, ?, ?]. Solutions [?]. Sorted [?]. Source [?, ?]. Sources [?, ?]. Spam [?, ?, ?]. Special [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Specific [?, ?]. Spell [?]. SPMDL [?]. Spread [?, ?]. Stance [?, ?]. standard [?]. standardized [?]. Standards [?, ?]. start [?]. Starting [?]. State [?]. State-of-the-art [?]. Statements [?]. Stemming [?]. Step [?]. Stock [?]. Stop [?]. Stop-and-Frisk [?]. Stopping [?, ?]. Stores [?]. Streaming [?, ?, ?, ?]. Strength [?]. String [?]. structured [?]. Student [?, ?]. Studies [?, ?]. Study [?, ?, ?, ?, ?, ?, ?, ?, ?]. Subjectivity [?]. Succeeding [?]. Success [?]. Supervised [?]. Support [?, ?, ?, ?]. Supporting [?]. Survey [?, ?, ?, ?, ?, ?, ?]. Swipe [?]. Syntax [?]. Synthesis [?]. Synthesize [?]. Synthetic [?, ?]. System [?, ?, ?, ?, ?, ?, ?, ?]. Systematic [?, ?, ?]. Systems [?, ?, ?, ?, ?, ?, ?].

t [?]. t-SNE [?]. Targeted [?]. Task [?]. Tasks [?, ?]. Taxonomy [?]. Teacher [?]. Teacher-Student [?]. Teaching [?]. Technical [?]. Technique [?, ?]. Techniques [?, ?, ?, ?]. Technologies [?]. Temporal [?, ?, ?]. TENSAI [?]. Term [?, ?]. Test [?, ?]. Text [?, ?, ?, ?]. Text2EL [?]. Textual [?, ?, ?]. Their [?]. theoretic [?]. Theory [?]. Things [?, ?, ?]. Thinking [?]. Time [?, ?, ?, ?, ?, ?, ?, ?, ?]. Time-Series [?]. tomography [?]. Tool [?]. Tools [?]. Topic [?]. Topics [?]. Toxicity [?]. Tradeoffs [?]. Traffic [?]. Training [?, ?]. Transactional [?]. Transfer [?]. Transformation [?]. Transformer [?]. Transformer-based [?]. Transforming [?]. Transparency [?, ?, ?, ?, ?, ?, ?, ?, ?]. Transparency-Data [?]. Transparency-Uses [?]. Transparent [?]. trial [?]. True [?]. Trust [?, ?, ?]. Truth [?, ?, ?]. Truthful [?]. Truthfulness [?, ?]. Twitter [?, ?, ?]. Types [?, ?]. Typing [?].

Ubiquitous [?]. Uncertain [?, ?]. Uncertainty [?, ?, ?]. Understand [?]. Understanding [?, ?]. Underwater [?]. Uniform [?]. Unifying [?]. Unintentional [?]. Union [?]. Uniqueness [?]. Unknown [?]. Unknowns [?]. Unlocking [?]. Unstructured [?]. Unsupervised [?, ?, ?]. Unverified [?]. Updating [?]. Urban [?]. Usage [?]. Use [?, ?, ?]. Used [?]. User [?, ?, ?, ?, ?, ?, ?]. User-Generated [?, ?, ?]. Users [?]. Uses [?]. Using [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Utilizing [?, ?].

Validating [?, ?]. Validation [?]. Validity [?]. Valuation [?, ?]. Value [?, ?]. Value-Based [?]. Values [?, ?, ?]. Various [?]. Vehicles [?]. Veracity [?, ?, ?]. versus [?]. Video [?, ?]. Virtual [?]. Vision [?, ?, ?]. Visits [?]. Visual [?]. Visualization [?, ?]. Volume [?]. Vulnerabilities [?].

Waste [?]. Web [?, ?, ?, ?, ?, ?, ?]. Well [?, ?]. Where [?]. Which [?, ?].

Why-Not [?]. Wikidata [?]. Within [?, ?, ?]. without [?, ?, ?]. Word [?]. Workflow [?]. Workflows [?]. Zaveri [?].